OpenStudy (luigi0210):

So I have no idea on how to do implicit differentiation:

OpenStudy (luigi0210):

Find y' \[\LARGE xy=x-2y\]

OpenStudy (anonymous):

What's the derivative of xy using implicit differentiation?

OpenStudy (luigi0210):

So I'll make an attempt: \[\LARGE \frac{d}{dx}(xy)=\frac{d}{dx}(x-2y)\] \[\LARGE (xy')+(y)=(1-2y')\] \[\LARGE (xy'+y)-(1-2y')=0\] Right so far?

OpenStudy (jbo11):

yea

OpenStudy (anonymous):

Just solve for y' now

OpenStudy (luigi0210):

So next is: \[\LARGE xy'+2y'=-y+1\] \[\LARGE y'(x+2)=-y+1\] and tada? \[\LARGE y'=-\frac{y+1}{x+2}\]

OpenStudy (anonymous):

Right

OpenStudy (luigi0210):

huh, well thank you c:

OpenStudy (anonymous):

No problem